A prism has a refracting angle of 60 degrees and a refractive index of 1.60. calculate the angle of incidence on the first face for a ray that is just totally reflected at the second face of the prism


1
Expert's answer
2021-10-10T16:02:17-0400
"\theta_c=\arcsin\frac{n_1}{n_2}=38.7°."

From the diagram, find the angle of refraction:


"\theta_r=60°-\theta_c=21.3°."

Apply Snell's law:


"\frac{\sin\theta}{\sin\theta_r}=\frac{n_2}{n_1},\\\space\\
\theta=\arcsin\bigg(\sin\theta_r·\frac{n_2}{n_1}\bigg)=35.5°."